Consider a study designed to compare different joystick types on the time it takes to complete a navigation mission.

A group of technology students is interested in whether haptic feedback (forces and vibrations applied through a joystick) is helpful in navigating a simulated game environment they created. To investigate this, they randomly assign 20 students to each of three joystick controller types and record the time it takes to complete a navigation mission. The joystick types are (1) a standard video game joystick, (2) a game joystick with force feedback, and (3) a game joystick with vibration feedback.

The three joystick types are designated 1, 2, and 3 and the parameters of the model are

μ1, μ2, μ3,

and σ. The following table summarizes the time (seconds) data.

Joystick

x

s n
1 279 72 20
2 245 62 20
3 258 82 20

(a)

Is it reasonable to pool the standard deviations for these data? Explain your answer.

Yes, the smallest s is half the largest s.

Yes, the largest s is less than twice the smallest s.    

No, the largest s is more than twice the smallest s.

No, the largest s is exactly twice the smallest s.

There is not enough information.

(b)

For each parameter,

μ1, μ2, μ3,

and σ, give the estimate. (Round your answer for σ to three decimal places.)

μ1=

μ2=

μ3=

σ=
Answer #1

a)

Yes, the largest s is less than twice the smallest s.    

b)

from above:

Group ni SS=(ni-1)*s2
1 20 98496.0000
2 20 73036.0000
3 20 127756.0000
grand mean= 299288.0000
SSE

μ1= 279

μ2= 245

μ3= 258

σ= sqrt(SSE/(N-k))=sqrt(299288/57) =5250.667